Title

INAF queue observing, 2014 Oct 29-30

Observer: E. Sani, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: M. Ewards on call
Instrument: LUCI, MODS, LBC-PARALLEL

Summary

When the Dome is opened the sky is partly cloudy, the humidity 40%, velocity of wind is 10 m/s After the transit thick clouds

Data Summary

Project_RUN Objects Filter ExpTime Notes
Nascimbeni HAT32 dual grating + red channel (LBC) 5 h 30m COMPLETED
standard g191-b2b dual grating    
Antoniucci dtauri dual grating + imaging   to be repeated
Antoniucci dtauri Ks   to be repeated
11.30 Mods flat fields slit 5 arcsec, dual grating with lamp QTH+ UG5 VF5 mods1b.0001-0003, modsr1.0001-0003

11.55 Mods flat fields slit 5 arcsec, red grating with lamp QTH+ UG5 VF10 mods1r.0004-0006

2.45 AcqMOds hat32.acq mods1r.20141031.0007-011

2.50 Collimations of LBC red channel
  • dofpia,/first,/X2,/redonly

03.15 Run defocus.ob a single 60s exposure to check the counts. the TO have done a defocus of -0.6 and the counts are ok.

03.22 ExecMods hat mods.hat32.obs mods1r.0012-96, mod1b.0001-58 and contemporary run hat32_f972.ob
  • average seeing is 1.1

* 3.55* defocus 0f -0.1 but the counts are good (25000 adu)
  • 04:10 clouds passing. See the attached plot

04.28 Seeing 1.6 on the GCS. And extremely variable (see the attached plot, N.B. DIMM measures higher values than the GCS)

07:55 defocus of +0.5

08:44 the seeing is 1.6-1.7, we slew to Antoniucci program
  • first we acquire the spectrum of the standard star g191-b2b with the slit 0.8'' acqMods g191b2b.acq mods1r.0097-0101
  • the slit is so narrow that modsAlign has not able to correct the offset, we don't see the standard in the slit. we try to insert a wider slit but also in this of 5'' the standard is out
  • we try the procedure with -r mods1r.0102-103

09:21 expecMODS g191b2b.ob mods1r.0104-0106,mods1b.0059-0061 * mods1r. 0108 spectrum of standard with slit 0.8'' exptime 60'' * mods1r. 0109-0111 slit images of standard with slit 0.8'' exptime 60''

09:95 expecMODS g191b2b.ob with slit 0.8'' mods1r.0112-114, mods1b.0062-064

10:09 slew to scientific target V512-Per, acqMODS v512per_ls.acq mods1r.0115-0116

10.25 since there are thick clouds and the guiding stars are faint, we have to switch to another targer

10.30 acqMODS drtau_ls.obs mods1r.0017-0019

10.42 execMODS dtrtau_ls.obs mods1r.0120-0123, mods1b.0064-0068==

11.02 execMODS drtau_ls.img. the guider has lost the guide star, we try again, but there is not collimation.
  • mods1r.0124-126,mods1b.0069-0071 there is'nt signal, it's too cloudy
  • we try again increases exp time to 60 sec mods1r.0127-129,mods1b.0072-0074

11.25 we switch to LUCI

11.50 we run the standard
  • executeLUCIscript.sh hip27881_ls_zJ.acq luci.0001-008
  • it is not possible to find the guide star, too cloudly.
  • the wind is increasing

12.05 we switch to the scientific target in imaging execLUCIscript.sh drtau_H.img luci.0009-0012
  • bad seeing , thick clouds passing.

-- %USERSIG{RobertaCarini - 2014-10-30}%

Comments

 
Topic revision: r10 - 19 Dec 2014, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback